### 08 Mar 2024 v1

# RadarDistill: Boosting Radar-based Object Detection Performance via Knowledge Distillation from LiDAR Features

Abstract
The inherent noisy and sparse characteristics of radar data pose challenges in finding effective representations for 3D object detection. In this paper, we propose RadarDistill, a novel knowledge distillation (KD) method, which can improve the representation of radar data by leveraging LiDAR data. RadarDistill successfully transfers desirable characteristics of LiDAR features into radar features using three key components: Cross-Modality Alignment (CMA), Activation-based Feature Distillation (AFD), and Proposal-based Feature Distillation (PFD). CMA enhances the density of radar features through multiple layers of dilation operations, effectively addressing the challenges of inefficient knowledge transfer from LiDAR to radar. AFD is designed to transfer knowledge from significant areas of the LiDAR features, specifically those regions where activation intensity exceeds a predetermined threshold. PFD guides the radar network to mimic LiDAR network features in the object proposals for accurately detected results while moderating features for misdetected proposals like false positives. Our comparative analyses conducted on the nuScenes datasets demonstrate that RadarDistill achieves state-of-the-art (SOTA) performance for radar-only object detection task, recording 20.5% in mAP and 43.7% in NDS. Also, RadarDistill significantly improves the performance of the camera-radar fusion model.
Figure 1: Illustration of RadarDistill. Through knowledge transfer from LiDAR features to radar features, our RadarDistill converts sparse and noisy radar features into dense and semantically rich LiDAR-like features for better object detection.
1 Introduction
While 3D perception based on camera and LiDAR sensors has been widely studied, radar sensors are now gaining attention due to their affordability and reliability in adverse weather conditions. Radar sensors can locate objects in a Bird’s Eye View (BEV) and also measure their radial velocity using Doppler frequency analysis. However, when compared to LiDAR or camera sensors, radar’s major limitations are its lower spatial resolution and a higher likelihood of false positives due to multi-path reflections. For decades, traditional object detection and tracking methods, based on manually crafted models, have been developed to overcome these limitations by many radar manufacturers. While deep neural networks (DNNs) have considerably improved 3D perception in camera and LiDAR sensors, similar advancements have not been mirrored in radar sensor-specific architectures. There are only a handful of studies that have applied deep neural networks to radar data. For instance, KPConvPillars [28] and Radar-PointGNN [26] leveraged KPConv [27] and graph neural networks, respectively, to detect objects using radar point clouds. However, these methods have not yet achieved the level of significant improvements realized with camera or LiDAR data. Recently, it was shown that radar can be effectively fused with camera or LiDAR data to enhance the robustness of 3D object detection [14, 15, 42, 21, 31, 13].
This paper focuses on improving the performance of radar-based 3D object detection using deep neural networks. We note that the limited performance of radar is largely due to the challenges in finding effective representations, given the sparse and noisy nature of radar measurements. Inspired by the remarkable success of deep models that encode LiDAR point clouds, our goal is to transfer the knowledge extracted from a LiDAR-based model to enhance a radar-based model. Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby refining the representation quality of the target model. To date, various KD methods have been introduced in the literature [8, 4, 11, 3, 41, 16, 32, 40]. Cross-modality knowledge distillation can be broadly divided into two approaches based on the modality of the student model used. When employing a camera as the student model, both depth and shape information are transferred from the teacher model to the student model. BEVDistill [3] transforms both LiDAR and cameras features into a BEV format, enabling the transfer of spatial knowledge from LiDAR features to camera features. DistillBEV [32] utilizes the prediction results from either LiDAR or LiDAR-camera fusion models to distinguish between foreground and background, guiding the student model to focus on KD in essential areas. S2M2-SSD [40] determines key areas based on the student model’s predictions and transfers information obtained from a LiDAR-camera fusion model in the key areas. Apart from these approaches, UniDistill [41] employs a universal cross-modality framework that enables knowledge transfer among diverse modalities. This framework is adaptable to different modality pairings, including camera-to-LiDAR, LiDAR-to-camera, and (camera+LiDAR)-to-camera settings.
In this paper, we propose RadarDistill, a novel KD framework designed to enhance the representation of radar data, leveraging LiDAR data. Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich features derived from LiDAR data for better object detection. Although both LiDAR data and its encoding network are used to enhance radar features during the training phase, they are not required in the inference phase.
Our proposed RadarDistill is designed based on three main ideas: 1) Cross-Modality Alignment (CMA), 2) Activation-based Feature Distillation (AFD), and 3) Proposal-based Feature Distillation (PFD). Our study indicates that transferring knowledge from LiDAR to radar features is difficult due to the inherent sparsity of radar data, which complicates the alignment with the more densely distributed LiDAR features. To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of activated radar features by implementing multiple layers of dilation operations.
The proposed AFD and PFD minimize the distribution gap between the intermediate features produced by the radar and LiDAR encoding networks. Initially, AFD conducts Activation-aware Feature Matching on low-level features. Specifically, it divides both radar and LiDAR features into active and inactive regions according to activation intensity of each features and constructs KD losses for each region separately. By assigning greater importance to the KD loss linked with active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
Next, PFD implements Proposal-level Selective Feature Matching, aimed at narrowing the differences from features associated with the proposals generated by the radar detection head. PFD directs the radar network to generate object features that are similar in shape to the high-level LiDAR features for accurately detected proposals. Conversely, for misdetected proposals such as false positives, the model is guided to suppress falsely activated features, reflecting low activation of the LiDAR features.
[…]
The key contributions of our work are as follows:
  • Our study is the first to demonstrate that radar object detection can be substantially improved using LiDAR data during the training process. Our qualitative results in Fig. 1 highlight that the radar features acquired through RadarDistill successfully mimic those of LiDAR, leading to enhanced object detection and localization.
  • Our findings reveal that the CMA is a crucial element of RadarDistill. In its absence, we observed a significant drop in performance enhancement. According to our ablation study, CMA plays a pivotal role in resolving inefficient knowledge transfer caused by the different densities of radar and LiDAR point clouds.
  • We propose two novel KD methods, AFD and PFD. These methods bridge the discrepancy between radar and LiDAR features, operating at two separate feature levels and utilizing KD losses specifically designed for each level.
  • RadarDistill achieves the state-of-the-art performance in the radar-only object detector category on the nuScenes benchmark. It also achieves a significant performance boost for camera-radar fusion scenarios.
2 Related Works
2.1 Radar-based 3D Object Detection
Radar-only 3D object detection models have employed backbone models adopted from various LiDAR-based detectors to suit the specific needs of radar data. Radar-PointGNN [26] utilized GNNs [24] to effectively extract features from sparse radar point clouds through a graph representation. KPConvPillars [28] introduced a hybrid architecture that combined grid-based and point-based approaches for radar-only 3D object detection. Recent studies have concentrated on detection techniques that combine radar with LiDAR or cameras [21, 14, 15, 42, 34, 31]. These methods aimed to complement the limitations of each sensor by using the data obtained from radars. RadarNet [34] employed a voxel-based early fusion and an attention-based late fusion approach to integrate radar and LiDAR data. RCM-Fusion [13] fused radar and camera data at both the feature level and instance level to fully utilize the potential of radar information. CRN [15] leveraged radar data to transform camera image features into a BEV view and then combined these transformed camera features with the radar BEV features using multi-modal deformable attention.
2.2 Knowledge Distillation
Knowledge distillation, as a strategy for model compression, transfers the information from a larger teacher model with greater capacity to a smaller student model. The student model mimics the teacher’s intermediate features [22], prediction logit [10], or activation boundary [9] to acquire knowledge from the teacher model. KD approaches have recently been extended from image classification to object detection [2, 17, 29, 7, 37, 35]. Defeat [7] utilized KD by decoupling features to foreground and background with ground truth. FGD [37] employed spatial and channel attention for 2D object detection, guiding the student model to focus on critical pixels and channels.
In 3D object detection, KD was applied to reduce model complexity in 2D detection [36, 19, 39], or to enhance detection performance through cross-modality knowledge distillation [8, 4, 11, 18, 40, 3, 41, 16, 32]. SparseKD [36] focused on KD for primary regions identified by the teacher model’s predictions, and the student model was initialized with the weights of the teacher to enhance feature extraction. MonoDistill [4] projected LiDAR point clouds onto the image plane and used a 2D CNN to develop an ‘image-version’ LiDAR-based model, effectively bridging the gap in feature representations between LiDAR and camera. BEVDistill [3] projected LiDAR and camera features into the BEV space, effectively transferring 3D spatial knowledge through dense feature distillation and sparse instance distillation. UniDistill [41] proposed a universal framework that was suitable for various teacher-student pairs. UniDistill projected both teacher and student features into BEVs and applied KD on the object features within each ground truth box.
As compared to the existing studies on cross-modality KD methods, our study focuses on developing KD specifically tailored for radar object detection, considering the sparse and noisy nature of radar data.
Figure 2: Overall architecture of RadarDistill. The input point clouds from each modality are initially processed through Pillar Encoding and Sparse 2D CNN to extract low-level features. CMA processes the low-level features of the radar to make them denser and align them with the features of LiDAR. AFD distinguishes between active and inactive regions in both radar and LiDAR features and minimizes their KD losses using different importance factors. Subsequently, proposal-level features are identified using the detection head, and PFD reduces the distribution gap between proposal features obtained from the radar and LiDAR networks. LiDAR data is only needed during the training phase to enhance radar features and not during inference.
3 Method
In this section, we present the details of the proposed KD framework, RadarDistll. Fig. 2 illustrates the overall architecture of RadarDistill. In Section 3.1, we provide a brief overview of the baseline model employed as both teacher and student models. In Section 3.2, we describe the details of Cross-Modality Alignment, a module for densifying radar features. In Section 3.3, we present the AFD, a novel Activation-aware Feature Matching approach. Finally Section 3.4 outlines the PFD, a Proposal-level Selective Feature Matching approach.
3.1 Preliminary
We used PillarNet [23] as our baseline model for both LiDAR and radar detectors. PillarNet initially generates regular pillars from the point clouds, subsequently producing sparse 2D pillar features denoted as $F^{2D}_{ldr}$ for LiDAR and $F^{2D}_{rdr}$ for radar through the pillar feature network. Then, we define the low-level BEV features denoted by $F^{l}_{ldr}$ and $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding 2D pillar features with a 2D sparse convolution-based encoder (SparseEnc), as follows
$F^{l}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
where low-level BEV features are sparse features derived from processing through SPConv [6] and $mod$ represents the detector modality, either ldr or rdr. We describe the high-level BEV features denoted by $F^{h}_{ldr}$ and $F^{h}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$, formed by encoding low-level BEV features $F^{l}_{ldr}$ and $F^{l}_{rdr}$ using a 2D dense convolution-based encoder (DenseEnc) as
$F^{h}_{mod}=DenseEnc(F^{l}_{mod}).$ (2)
The high-level BEV features are processed through a centerpoint[38]-based head (CenterHead) to generate the final prediction heatmaps for classification, regression, and IoU scoring, represented as $H^{cls}$, $H^{reg}$ and $H^{IoU}$. The final predictions are obtained from
$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{h}_{mod}).$ (3)
3.2 Cross-Modality Alignment
Our goal is to reduce the disparity between radar and LiDAR due to the differences in point cloud density via CMA. PillarNet [23] employs SPConv [6] to generate low-level features from only non-empty pillars. Notably, when comparing the number of non-empty pillars, radar’s non-empty pillars ($F^{l}_{rdr}$) comprise only $11\%$ of the total count of LiDAR’s non-empty pillars ($F^{l}_{ldr}$). This significant disparity in the quantity of non-empty pillars necessitates an alignment of the two modalities, especially since the information from LiDAR’s non-empty pillars cannot be directly transferred to the corresponding empty pillars in radar data.
Fig. 3 presents the structure of the CMA. Within the component of CMA, the down block applies deformable convolution [5] to selectively extract essential features during downsampling, and then it aggregates these features using the ConvNeXt V2 blocks [33]. In the up block, dilation is applied using 2D transposed convolution to densify the features in surrounding areas. The aggregation module that includes concatenation followed by a 1$\times$1 convolution layer combines different features and operates similarly to a skip connection.
We downsample the 1/8 resolution features $F^{l}_{rdr}$ to produce $F^{l^{\prime(1)}}_{rdr}$. These $F^{l^{\prime(1)}}_{rdr}$ features are upsampled and combined with $F^{l}_{rdr}$ in the aggregation module, forming the aligned BEV features $F^{a^{(1)}}$. Subsequently, we reduce the resolution of $F^{l^{\prime(1)}}_{rdr}$ to 1/32 and then increase it to 1/16. These upsampled features are integrated with the downsampled $F^{a^{(1)}}$, to produce $F^{l^{\prime(2)}}_{rdr}$. Finally, we upsample the $F^{l^{\prime(2)}}_{rdr}$ and fuse it with $F^{a^{(1)}}$ in the aggregation module, which generates the second aligned BEV features $F^{a^{(2)}}$. In the CMA process, we acquire two sets of aligned features at 1/8 resolution. We apply multiple AFD to these aligned features to enhance their similarity to LiDAR’s low-level features. Consequently, as observed in Fig. 1, the radar’s low-level features become denser after being processed through the CMA. A comprehensive description of the CMA’s component is available in the Supplementary Materials.
3.3 Activation-based Feature Distillation
AFD strategically aligns the low-level features of radar and LiDAR through Activation-aware Feature Matching. This process matches the activation patterns of radar features to those of LiDAR, thus bridging the gap in their distribution of active features. Additionally, AFD is applied iteratively to each pair of the $n$-th aligned features, $F^{a^{(n)}}$ and $F^{l}_{ldr}$, ensuring a more precise and effective alignment between the radar and LiDAR features. First, for both LiDAR and radar features, we define “active regions” as areas where each radar and LiDAR feature intensity values, summed in the channel direction, are greater than 0. Conversely, areas where these sums are less than 0 are classified as “inactive regions”. We define active mask $M^{mod}_{i,j}$ to present these regions and obtain the active masks $M^{a^{(n)}}_{i,j}$and $M^{ldr}_{i,j}$ for $F^{a^{(n)}}$and $F^{l}_{ldr}$ correspondingly. These masks are obtained from
$F^{mod}_{i,j}=\sum^{C}_{c=1}F^{mod}_{c,i,j},$ (4)
$M^{mod}_{i,j}=\begin{cases}1,&if\,(F^{mod}_{i,j}>0),\\ 0,&otherwise,\end{cases}$ (5)
where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, the indices i,j,c correspond to height, width, and channel, and $mod$ indicates the modality of features, comprising either ldr or a${}^{(n)}$. The distillation region is adaptively separated into two categories: the active regions (AR) where both radar and LiDAR are active and the inactive regions (IR) where radar is active while LiDAR is inactive. There is an imbalance between the AR and IR due to the sparse and noisy characteristics of radar. This can disturb training by focusing on imitating the dominant IR. Therefore, we apply a relative adaptive weight based on the number of pixels in each region. AR is assigned a weight of 1 as it is a smaller yet crucial region while IR is given a relative adaptive weight denoted by $\rho$ as determined by the ratio of pixel counts between AR and IR. We define a separate mask $M^{sep^{(n)}}$ for distinguishing the regions, specifically between $F^{a^{(n)}}$and $F^{l}_{ldr}$ as
$AR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=1),$ (6)
$IR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=0),$ (7)
$M^{sep^{(n)}}_{i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(n)},\\ \rho^{(n)}\times\beta,&if\,(i,j)\in IR^{(n)},\\ 0,&\,otherwise,\end{cases}$ (8)
$\rho^{(n)}$ $=\frac{N_{AR^{(n)}}}{N_{IR^{(n)}}},$ (9)
where $\rho^{(n)}$ represents the relative importance of the $IR^{(n)}$ in comparison to $AR^{(n)}$, $\alpha$ and $\beta$ are the balancing parameters between the separated regions, and $N$ is the number of pixels. In this way, the distillation region and the $\rho^{(n)}$ value are adaptively adjusted for each training, and distillation is performed with different weights for AR and IR.
Figure 3: Detailed structure of the proposed CMA module.
After decomposing the distillation region with AR and IR, the $n$-th aligned features $F^{a^{(n)}}_{rdr}$ are trained to imitate $F^{l}_{ldr}$. The distillation loss $L^{(n)}_{low}$ for the $n$-th activation features is given by
$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{sep^{(n)}}_{i,j}(F^{ldr}_{c,i,j}-F^{a^{(n)}}_{c,i,j})^{2},$ (10)
where $F^{ldr}$ denotes the LiDAR’s low-level BEV features $F^{l}_{ldr}$.
The active mask loss is applied between the predicted mask, obtained by applying a sigmoid function to $F^{a(n)}$, and $M^{ldr}$ to ensure the radar’s active mask closely mimics LiDAR. The active mask loss $L^{(n)}_{mask}$ is calculated as
$L^{(n)}_{mask}$ $=\sum^{H}_{i=1}\sum^{W}_{j=1}|M^{ldr}_{i,j}-\sigma(F^{a^{(n)}}_{i,j})|,$ (11)
where $\sigma(\cdot)$ denotes a sigmoid function.
We apply AFD twice to the aligned BEV features to more accurately imitate $F^{l}_{ldr}$, and use the average of losses calculated from these multiple AFD as the final AFD loss. The final AFD loss function is given by
$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}(L^{(n)}_{low}+L^{(n)}_{mask}).$ (12)
3.4 Proposal-based Feature Distillation
PFD employs Proposal-level Feature Matching to reduce the discrepancies between the high-level features. It utilizes predictions generated from the radar, guiding the radar network to create object features that closely mimic the high-level features of LiDAR, thereby achieving a more accurate alignment in properly identified proposals. Additionally, for erroneously identified proposals, such as false positives, PFD directs the model to suppress falsely activated features, thereby mirroring the lower activation levels typically seen in LiDAR features for such instances.
We employ the radar’s predicted heatmap to adaptively define the crucial region according to radar. Additionally, there is a significant scale disparity between the high-level BEV features generated by radar and LiDAR. PFD aligns the scales between the two modalities by making the student mimic the normalized values that are applied in the channel direction for both radar and LiDAR. DenseEnc perform input $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ by first downsampling to 1/16 of its size and then upsampling to 1/8 to generate the initial high-level BEV feature $F^{h^{(1)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Subsequently, $F^{h^{(1)}}_{rdr}$ and CMA’s output $F^{a^{(2)}}_{rdr}$ are concatenated and then fed through a convolution layer resulting in the second high-level BEV feature $F^{h^{(2)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$. Similar to AFD, PFD is applied twice to each of the multiple feature sets $F^{h^{(m)}}_{rdr}$.
The first step involves thresholding the predicted radar heatmap $H^{cls}_{rdr}$ and the GT heatmap $H^{cls}_{GT}$ to identify the regions of true positives, false positives, and false negatives. The regions are identified using the following operations
$TP$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}>\sigma),$ (13)
$FP$ $=\,(H^{cls}_{GT}<\sigma)\,\&\,(H^{cls}_{rdr}>\sigma),$ (14)
$FN$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}<\sigma),$ (15)
where $H^{cls}_{rdr}$ represents the predicted radar classification heatmap, $H^{cls}_{GT}$ is the GT classification heatmap, and $\sigma$ is the threshold parameter set at 0.1.
TP and FN are considered important regions according to the foreground, while FP indicates regions where radar is mostly wrong. Consequently, TP and FN are categorized into one group, FP into another, leading to the generation of a scaling mask $M^{scale}$ for each defined area. The scaling mask is expressed as
$M^{scale}_{i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (16)
where $N$ denotes the number of pixels, while $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters for the respective regions. The alignment of feature scales between radar and LiDAR involves applying the channel-wise softmax to the high-level BEV features $F^{h}_{ldr}$ and $F^{h^{(m)}}_{rdr}$, which results in $S^{ldr}$ and $S^{h^{(m)}}$, i.e.,
$S^{mod}_{c,:,:}=\frac{{\text{e}xp}(\frac{F^{mod}_{c,:,:}}{\tau})}{\sum^{C}_{c^{*}=1}{\text{e}xp}(\frac{F^{mod}_{c^{*},:,:}}{\tau})},$ (17)
where $mod\in\{ldr,h^{(m)}\}$, $\tau$ is a hyper-parameter that controls the smoothing of the distribution and is used as 0.5. Similar to AFD, PFD is employed twice to train $S^{h^{(m)}}$ to imitate $S^{ldr}$ and is computed as
$L^{(m)}_{high}=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{scale}_{i,j}|S^{ldr}_{c,i,j}-S^{h^{(m)}}_{c,i,j}|.$ (18)
The final PFD loss value is derived from the average of the PFD loss values computed from multiple PFD. The final PFD loss is calculated from
$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}(L^{(m)}_{high}).$ (19)
3.5 Overall Loss Function
We freeze the teacher model that has been pre-trained on LiDAR point clouds and then train the student model end-to-end using radar point clouds. Finally, we compute the total loss as the weighted sum of the detection loss and the distillation losses as
$L_{total}=L_{det}+\gamma L_{AFD}+\delta L_{PFD},$ (20)
where $\gamma$ and $\delta$ are hyperparameters used to balance the scale of different distillation losses.
4 Experiments
4.1 Experimental Setup
Datasets and metrics. We conduct the experiments on the nuScenes [1] dataset. This dataset contains 700, 150, 150 driving scenes for training, validation, and testing each captured with images from six cameras, points from five radars and one LiDAR. Our evaluation utilizes the official metrics from nuScenes, which include mean average precision (mAP) and the nuScenes detection score (NDS).
[…]
Radar-PointGNN [26] R - 0.5 3.4 1.024 0.859 0.897 1.020 0.931
KPConvPillars [28] R - 4.9 13.9 0.823 0.428 0.607 2.081 1.000
PillarNet* [23] (Our baseline) R - 8.6 34.7 0.532 0.283 0.615 0.438 0.092
  RadarDistill R 20.5 43.7 0.461 0.263 0.525 0.336 0.072
Table 1: Performance comparison on nuScenes test set. The best performed metrics in the radar-only model are bolded. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves the state-of-the-art performance in all metrics.
Method Input KD Car Truck bus Trailer C.V Ped. Motor. Bicycle T.C Barrier
CenterFusion [21] C, R - 50.9 25.8 23.4 23.5 7.7 37.0 31.4 20.1 57.5 48.4
FCOS3D [30] C - 52.4 27.0 27.7 25.5 11.7 39.7 34.5 29.8 55.7 53.8
MonoDIS [25] C - 47.8 22.0 18.8 17.6 7.4 37.0 29.0 24.5 48.7 51.1
CenterNet [43] C - 53.6 27.0 24.8 25.1 8.6 37.5 29.1 20.7 58.3 53.3
PillarNet* [23] (Our baseline) R - 41.8 11.6 8.4 6.5 0.0 7.4 1.0 0.0 1.1 8.1
  RadarDistill R 54.0 15.3 11.3 29.5 5.5 9.2 15.3 0.9 21.7 42.3
Table 2: Per-class comparison on nuScenes test set. The best performed metrics are bolded. ’C.V’, ’Ped.’, ’Motor.’, and ’T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
Implementation details. As a baseline model, we utilized PillarNet-18, which is PillarNet [23] with a ResNet-18 backbone. We used the Adam optimizer with a learning rate of 0.001, applying a one-cycle learning rate policy. We set weight decay at 0.01, and adjusting momentum between 0.85 and 0.95. We trained our baseline models on 4 NVIDIA RTX 3090 GPUs for 20 epochs, employing a total batch size of 16 with the Class-Balanced Grouping and Sampling (CBGS) strategy [44]. We applied random scene flipping along the axes, random rotation, scaling, translation and ground-truth box sampling as the data augmentation. We set our detection range on the $X$ and $Y$ axes to $[-54.0m,54.0m]$ and on the Z-axis to $[-3m,5m]$ , utilizing pillars of $(0.075m,0.075m)$. Our proposed model was trained for $40$ epochs and all other training procedures were applied identically to the baseline. We adopted the teacher’s weights to initialize the student model, following the inheriting strategy [12]. In Eq. (9), we set the hyper-parameters $\alpha$ to 3e-4 and $\beta$ to 5e-5, while in Eq. (16), the hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ were set to $5$ and $1$, respectively. Additionally, in Eq. (20), we set the parameter $\gamma$ and $\delta$ to $5$ and $25$.
4.2 Comparison with the State-of-the-Arts
We evaluated the performance of RadarDistill on nuScenes test set. In Table 1, our RadarDistill achieved top-ranking performance on the nuScenes radar-only benchmark, significantly outperforming other methods. Specifically, our model achieved a $20.5\%$ mAP and $43.7\%$ NDS, surpassing the previous state-of-the-art (SOTA) method, KPConvPillars [28], by a substantial margin of $+15.6\%$ in mAP and $+29.8\%$ in NDS.
As shown in Table 2, we compared the performance of our RadarDistill for each class with other camera-only and camera-radar fusion models. Our model outperformed the baseline model (PillarNet [23]), especially for the barrier and trailer classes, with gains of at least $+23\%$. Furthermore, our model demonstrated the best performance for the car and trailer classes even when compared with other camera-only models and camera-radar fusion model approaches. However, the pedestrian and bicycle classes showed only slight performance improvements, largely due to the inherent limitations of radar sensors, as these objects are smaller and have few radar points in their ground truth (GT) boxes.
4.3 Ablation Studies
We conducted several ablation studies using the nuScenes validation set. We trained our model on 1/7 of training set and used the complete validation set for evaluation to reduce the time spent on experiments.
Component Analysis. We present an ablation study to evaluate the impact of each component of our proposed model, as shown in Table 3. In #2, when CMA is applied to the baseline, there’s an overall enhancement in performance; yet, the improvements are relatively low. Comparing #2 with #3 and #4, where CMA is combined with AFD and PFD respectively, we see a significant performance gain. For #3, there is an increase in mAP by $+4.5\%$ and NDS by $+4.4\%$. In the case of #4, mAP improves by $+1.8\%$ and NDS by $+3.0\%$. Additionally, when all components are used together as in #6, we can see that it achieves the best performance across all metrics. Lastly, to understand the influence of CMA in the distillation process, #5 was tested without it. The results show a significant difference in performance, with an increase of $+4.2\%$ in mAP and $+4.0\%$ in NDS when CMA is included, demonstrating that the CMA structure performs a critical role in our model.
Comparative Analysis of Distillation Regions. We conducted experiments to analyze the importance of distillation regions during the knowledge distillation (KD) process.
Table 4 presents an ablation study conducted to evaluate the impact of selecting different regions when applying low-level feature distillation. When distillation is performed on the entire area without separate regions, we observe a performance improvement of $+4.2\%$ in Car AP, $+2.2\%$ in mAP, and $+2.9\%$ in NDS. Applying distillation to the gaussian heatmap areas centered around the GT boxes led to an increase in Car AP and mAP compared to the total, but a decrease in NDS. Dividing the distillation between foreground and background regions, compared to the total, resulted in an increase in both Car AP and mAP, but a slight decrease in NDS. Conversely, dividing the feature distillation by our proposed activation regions yielded the best performance across all metrics, with improvements of $+6.8\%$ in Car AP, $+2.3\%$ in mAP, and $+1.5\%$ in NDS compared to total.
Next, Table 5 presents an experiment comparing the performance between previous cross-modality KD methods and PFD’s approach in applying high-level feature distillation. Here, the first row is based on the application of AFD according to our proposed method. When distillation is performed on the entire area, we observe a slight decrease in Car AP and mAP by $-0.6\%$ and $-0.1\%$ respectively, while NDS shows a marginal increase of $0.1\%$, indicating a minimal change in performance. Even when comparing the gaussian heatmap areas centered around the GT boxes, only Car AP shows a slight increase $0.1\%$ with the rest being identical to the total. Furthermore, applying KD based on LiDAR prediction results also did not yield any notable performance improvements. In contrast, our method, which applies KD based on radar prediction boxes, showed an improvement of over $0.4\%$ across all metrics compared to the total, demonstrating the best performance.
Scale Alignment in PFD. We conducted experiments to examine the impact of aligning the scale between teacher and student in PFD process in Table 6. When applying PFD without normalization, we observe a substantial decrease of $-5.9\%$ in Car AP, $-2.4\%$ in mAP, and $-1.9\%$ in NDS compared to before applying PFD. Conversely, after implementing normalization and PFD, we observe an increase of $+0.3\%$ in Car AP, $+0.3\%$ in mAP, and $+1.0\%$ in NDS.
Apply radar-camera fusion model. Table 7 shows a comparison of the performance when using radar features enhanced by our method in the radar-camera fusion model. The enhanced radar and camera features were fused using the BEVFusion [20] method, with its detailed structure and training method provided in the Supplementary Materials. We observe a significant increase of $+1.8\%$ in Car AP, $+1.3\%$ in mAP, and $+1.1\%$ in NDS.
[…]
1 5.4 27.3
2 6.4 29.3
3 10.9 33.7
4 8.2 32.3
5 7.0 30.7
6 11.2 34.7
Table 3: Ablation study of components in Radardistill. “CMA”: Cross-Modality Alignment. “AFD”: Activation-based Feature Distillation. “PFD”: Proposal-based Feature Distillation.
Region AFD
[…]
W/O AFD 34.8 6.4 29.3
Total 39.0 8.6 32.2
GT 40.4 8.8 29.5
FG / BG 42.1 9.6 31.7
Activation 45.8 10.9 33.7
Table 4: Ablation study of distillation regions in AFD. “W/O AFD”: Applying only CMA without distillation. “FG/BG”: Foreground and background region.
Region PFD
[…]
W/O PFD 45.8 10.9 33.7
Total 45.2 10.8 33.8
GT 45.9 10.8 33.8
LiDAR prediction 45.4 10.8 34.0
Radar prediction 46.1 11.2 34.7
Table 5: Ablation study of distillation regions in PFD. “W/O PFD”: Applying only CMA and AFD. “LiDAR prediction” : Using only LiDAR prediction results. “Radar prediction” : Using only Radar prediction results
[…]
W/O PFD 45.8 10.9 33.7
W/O Norm 39.9 8.5 31.8
Norm 46.1 11.2 34.7
Table 6: Ablation study of scale alignment in PFD. “W/O PFD”: Applying only CMA and AFD. “W/O Norm”: Applying PFD without normalizing. “Norm”: Applying normalizing and PFD.
[…]
BEVFusion* [20] C,R 65.9 38.3 45.3
RadarDistill-CR C,R 67.7 39.6 46.4
Table 7: Ablation study of applying to a radar-camera fusion model. ”*” denotes models we have reproduced without applying test time augmentation. Our model achieves slight performance enhancements in all metrics.
5 Conclusion
In our study, we introduced a new radar-based 3D object detection method, RadarDistill, which leveraged LiDAR data via knowledge distillation for enhanced performance. Our approach successfully guided the radar encoding network to generate radar features that closely resemble the detailed and semantically rich features of LiDAR. CMA enhanced the student network’s expressive power beyond that of the teacher network, allowing the student network to learn the complex distribution of LiDAR features more effectively. Through Activation-aware Feature Matching and Proposal-level Selective Feature Matching, both AFD and PFD successfully reduced the discrepancy from LiDAR feature distributions, particularly in important regions within the intermediate features. Our experiments demonstrated that RadarDistill achieved the state-of-the-art performance in radar-only object detection and offered significant performance gains in camera-radar fusion scenarios. We leave the design of more sophisticated KD framework specifically tailored for camera-radar fusion for future work.

### 05 Apr 2024 v2

@@
 RadarDistill successfully transfers desirable characteristics of LiDAR features into radar features using three key components: Cross-Modality Alignment (CMA), Activation-based Feature Distillation (AFD), and Proposal-based Feature Distillation (PFD).
-CMA enhances the density of radar features through multiple layers of dilation operations, effectively addressing the challenges of inefficient knowledge transfer from LiDAR to radar.
-AFD is designed to transfer knowledge from significant areas of the LiDAR features, specifically those regions where activation intensity exceeds a predetermined threshold.
-PFD guides the radar network to mimic LiDAR network features in the object proposals for accurately detected results while moderating features for misdetected proposals like false positives.
+CMA enhances the density of radar features by employing multiple layers of dilation operations, effectively addressing the challenge of inefficient knowledge transfer from LiDAR to radar.
+AFD selectively transfers knowledge based on regions of the LiDAR features, with a specific focus on areas where activation intensity exceeds a predefined threshold.
+PFD similarly guides the radar network to selectively mimic features from the LiDAR network within the object proposals.
 Our comparative analyses conducted on the nuScenes datasets demonstrate that RadarDistill achieves state-of-the-art (SOTA) performance for radar-only object detection task, recording 20.5% in mAP and 43.7% in NDS.
 Also, RadarDistill significantly improves the performance of the camera-radar fusion model.
-Figure 1: Illustration of RadarDistill.
-Through knowledge transfer from LiDAR features to radar features, our RadarDistill converts sparse and noisy radar features into dense and semantically rich LiDAR-like features for better object detection.
+Figure 1: Illustration of Proposed RadarDistill.
+Our RadarDistill method facilitates knowledge transfer from LiDAR features to radar features, enhancing the quality of radar features for Bird’s Eye View (BEV) object detection.
 1 Introduction
@@
 Inspired by the remarkable success of deep models that encode LiDAR point clouds, our goal is to transfer the knowledge extracted from a LiDAR-based model to enhance a radar-based model.
-Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby refining the representation quality of the target model.
+Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby improving the representation quality of the target model.
 To date, various KD methods have been introduced in the literature [8, 4, 11, 3, 42, 16, 33, 41].
-Cross-modality knowledge distillation can be broadly divided into two approaches based on the modality of the student model used.
-When employing a camera as the student model, both depth and shape information are transferred from the teacher model to the student model.
 BEVDistill [3] transforms both LiDAR and cameras features into a BEV format, enabling the transfer of spatial knowledge from LiDAR features to camera features.
@@
 In this paper, we propose RadarDistill, a novel KD framework designed to enhance the representation of radar data, leveraging LiDAR data.
-Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich features derived from LiDAR data for better object detection.
+Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich LiDAR features.
 Although both LiDAR data and its encoding network are used to enhance radar features during the training phase, they are not required in the inference phase.
@@
 Our study indicates that transferring knowledge from LiDAR to radar features is difficult due to the inherent sparsity of radar data, which complicates the alignment with the more densely distributed LiDAR features.
-To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of activated radar features by implementing multiple layers of dilation operations.
+To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of active radar features by implementing multiple layers of dilation operations.
 The proposed AFD and PFD minimize the distribution gap between the intermediate features produced by the radar and LiDAR encoding networks.
@@
 Specifically, it divides both radar and LiDAR features into active and inactive regions according to activation intensity of each features and constructs KD losses for each region separately.
-By assigning greater importance to the KD loss linked with active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
+By assigning greater importance to the KD loss linked with sparsely distributed active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
 Next, PFD implements Proposal-level Selective Feature Matching, aimed at narrowing the differences from features associated with the proposals generated by the radar detection head.
@@
 It also achieves a significant performance boost for camera-radar fusion scenarios.
+Figure 2: Overall architecture of RadarDistill.
+The input point clouds from each modality are independently processed through Pillar Encoding followed by SparseEnc to extract low-level BEV features.
+CMA is then employed to densify the low-level BEV features in the radar branch.
+AFD then identifies active and inactive regions based on both radar and LiDAR features and minimizes their associated distillation losses.
+Subsequently, PFD conducts knowledge distillation based on proposal-level features obtained from DenseEnc.
+Note that the LiDAR branch is solely utilized during the training phase to enhance the radar pipeline and is not required during inference.
 2 Related Works
@@
 As compared to the existing studies on cross-modality KD methods, our study focuses on developing KD specifically tailored for radar object detection, considering the sparse and noisy nature of radar data.
-Figure 2: Overall architecture of RadarDistill.
-The input point clouds from each modality are initially processed through Pillar Encoding and Sparse 2D CNN to extract low-level features.
-CMA processes the low-level features of the radar to make them denser and align them with the features of LiDAR.
-AFD distinguishes between active and inactive regions in both radar and LiDAR features and minimizes their KD losses using different importance factors.
-Subsequently, proposal-level features are identified using the detection head, and PFD reduces the distribution gap between proposal features obtained from the radar and LiDAR networks.
-LiDAR data is only needed during the training phase to enhance radar features and not during inference.
 3 Method
@@
 3.1 Preliminary
-We used PillarNet [23] as our baseline model for both LiDAR and radar detectors.
-PillarNet initially generates regular pillars from the point clouds, subsequently producing sparse 2D pillar features denoted as $F^{2D}_{ldr}$ for LiDAR and $F^{2D}_{rdr}$ for radar through the pillar feature network.
-Then, we define the low-level BEV features denoted by $F^{l}_{ldr}$ and $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding 2D pillar features with a 2D sparse convolution-based encoder (SparseEnc), as follows
-$F^{l}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
-where low-level BEV features are sparse features derived from processing through SPConv [6] and $mod$ represents the detector modality, either ldr or rdr.
-We describe the high-level BEV features denoted by $F^{h}_{ldr}$ and $F^{h}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$, formed by encoding low-level BEV features $F^{l}_{ldr}$ and $F^{l}_{rdr}$ using a 2D dense convolution-based encoder (DenseEnc) as
-$F^{h}_{mod}=DenseEnc(F^{l}_{mod}).$ (2)
-The high-level BEV features are processed through a centerpoint[38]-based head (CenterHead) to generate the final prediction heatmaps for classification, regression, and IoU scoring, represented as $H^{cls}$, $H^{reg}$ and $H^{IoU}$.
-The final predictions are obtained from
-$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{h}_{mod}).$ (3)
+We use PillarNet [23] as our baseline model for both LiDAR and radar detectors.
+PillarNet organizes both radar and LiDAR point clouds using a 2D pillar structure of the same size.
+It then generates two BEV pillar features $F^{2D}_{ldr}$ and $F^{2D}_{rdr}$, utilizing separate pillar encoding networks in the LiDAR and radar branches, respectively.
+We define the low-level BEV features $F^{(l)}_{ldr}$ and $F^{(l)}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding the pillar features $F^{2D}_{ldr}$ and $F^{2D}_{rdr}$ as
+$F^{(l)}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
+where $SparseEnc(\cdot)$ denotes a 2D sparse convolution-based encoder (SparseEnc) [6] and $mod$ represents the detector modality, i.e., $mod\in\{ldr,rdr\}$.
+Similarly, we describe the high-level BEV features $F^{(h)}_{ldr}$ and $F^{(h)}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ formed by encoding the low-level BEV features $F^{(l)}_{ldr}$ and $F^{(l)}_{rdr}$, respectively as
+$F^{(h)}_{mod}=DenseEnc(F^{(l)}_{mod}),$ (2)
+where $DenseEnc(\cdot)$ denotes a 2D dense convolution-based encoder (DenseEnc).
+The high-level BEV features are further processed through a CenterHead network [39] to generate the final prediction heatmaps for classification, regression, and IoU scoring
+$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{(h)}_{mod}).$ (3)
+Note that the pipeline in the LiDAR branch serves as the teacher model, while the one in the radar branch acts as the student model in our KD framework.
+These networks, serving as teacher and student, respectively, do not share weights.
 3.2 Cross-Modality Alignment
-Our goal is to reduce the disparity between radar and LiDAR due to the differences in point cloud density via CMA.
-PillarNet [23] employs SPConv [6] to generate low-level features from only non-empty pillars.
-Notably, when comparing the number of non-empty pillars, radar’s non-empty pillars ($F^{l}_{rdr}$) comprise only $11\%$ of the total count of LiDAR’s non-empty pillars ($F^{l}_{ldr}$).
-This significant disparity in the quantity of non-empty pillars necessitates an alignment of the two modalities, especially since the information from LiDAR’s non-empty pillars cannot be directly transferred to the corresponding empty pillars in radar data.
+The objective of CMA is to enhance the density of radar BEV features, thereby facilitating the transfer of knowledge from LiDAR features to radar features more effectively.
+We note that the average number of non-empty pillars in radar features $F^{(l)}_{rdr}$ is only approximately 10% of the average number found in LiDAR features $F^{(l)}_{ldr}$.
+This considerable difference in the number of non-empty pillars poses challenges in aligning features between the two modalities.
+Specifically, transferring information from non-empty pillars in LiDAR data to corresponding empty pillars in radar data proves to be impractical.
+To tackle this challenge, CMA is employed to densify radar features.
+Our empirical study demonstrates that this densification process significantly contributes to the successful knowledge distillation from LiDAR to radar.
 Fig.
-3 presents the structure of the CMA.
-Within the component of CMA, the down block applies deformable convolution [5] to selectively extract essential features during downsampling, and then it aggregates these features using the ConvNeXt V2 blocks [33].
-In the up block, dilation is applied using 2D transposed convolution to densify the features in surrounding areas.
-The aggregation module that includes concatenation followed by a 1$\times$1 convolution layer combines different features and operates similarly to a skip connection.
-We downsample the 1/8 resolution features $F^{l}_{rdr}$ to produce $F^{l^{\prime(1)}}_{rdr}$.
-These $F^{l^{\prime(1)}}_{rdr}$ features are upsampled and combined with $F^{l}_{rdr}$ in the aggregation module, forming the aligned BEV features $F^{a^{(1)}}$.
-Subsequently, we reduce the resolution of $F^{l^{\prime(1)}}_{rdr}$ to 1/32 and then increase it to 1/16.
-These upsampled features are integrated with the downsampled $F^{a^{(1)}}$, to produce $F^{l^{\prime(2)}}_{rdr}$.
-Finally, we upsample the $F^{l^{\prime(2)}}_{rdr}$ and fuse it with $F^{a^{(1)}}$ in the aggregation module, which generates the second aligned BEV features $F^{a^{(2)}}$.
-In the CMA process, we acquire two sets of aligned features at 1/8 resolution.
-We apply multiple AFD to these aligned features to enhance their similarity to LiDAR’s low-level features.
-Consequently, as observed in Fig.
-1, the radar’s low-level features become denser after being processed through the CMA.
-A comprehensive description of the CMA’s component is available in the Supplementary Materials.
+3 illustrates the architecture of CMA.
+CMA comprises the Down Block, Up Block, and Aggregation Module.
+Down Block conducts down-sampling via deformable convolution [5], followed by the ConvNeXt V2 blocks [34].
+The Up Block conducts up-sampling operations using a 2D transposed convolution.
+The Aggregation Module concatenates two input features and applies a 1$\times$1 convolution layer.
+After applying the Down Block and Up Block operations twice each, in addition to incorporating a side pathway through the Aggregation Module, CMA generates two low-level BEV features, $F^{(l_{1})}_{rdr}$ and $F^{(l_{2})}_{rdr}$, at a $1/8$ resolution.
+As a result of CMA’s dilation operation, these low-level features exhibit increased density compared to the input radar features.
+(Refer to Fig.
+2 for a visual comparison of the features before and after CMA.) These intermediate features are subsequently utilized for knowledge distillation in the following AFD stage.
+For a detailed structure of the CMA, refer to the Supplementary Materials.
+Figure 3: Detailed structure of the proposed CMA module.
 3.3 Activation-based Feature Distillation
-AFD strategically aligns the low-level features of radar and LiDAR through Activation-aware Feature Matching.
-This process matches the activation patterns of radar features to those of LiDAR, thus bridging the gap in their distribution of active features.
-Additionally, AFD is applied iteratively to each pair of the $n$-th aligned features, $F^{a^{(n)}}$ and $F^{l}_{ldr}$, ensuring a more precise and effective alignment between the radar and LiDAR features.
-First, for both LiDAR and radar features, we define “active regions” as areas where each radar and LiDAR feature intensity values, summed in the channel direction, are greater than 0.
-Conversely, areas where these sums are less than 0 are classified as “inactive regions”.
-We define active mask $M^{mod}_{i,j}$ to present these regions and obtain the active masks $M^{a^{(n)}}_{i,j}$and $M^{ldr}_{i,j}$ for $F^{a^{(n)}}$and $F^{l}_{ldr}$ correspondingly.
-These masks are obtained from
-$F^{mod}_{i,j}=\sum^{C}_{c=1}F^{mod}_{c,i,j},$ (4)
-$M^{mod}_{i,j}=\begin{cases}1,&if\,(F^{mod}_{i,j}>0),\\ 0,&otherwise,\end{cases}$ (5)
-where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, the indices i,j,c correspond to height, width, and channel, and $mod$ indicates the modality of features, comprising either ldr or a${}^{(n)}$.
-The distillation region is adaptively separated into two categories: the active regions (AR) where both radar and LiDAR are active and the inactive regions (IR) where radar is active while LiDAR is inactive.
-There is an imbalance between the AR and IR due to the sparse and noisy characteristics of radar.
-This can disturb training by focusing on imitating the dominant IR.
-Therefore, we apply a relative adaptive weight based on the number of pixels in each region.
-AR is assigned a weight of 1 as it is a smaller yet crucial region while IR is given a relative adaptive weight denoted by $\rho$ as determined by the ratio of pixel counts between AR and IR.
-We define a separate mask $M^{sep^{(n)}}$ for distinguishing the regions, specifically between $F^{a^{(n)}}$and $F^{l}_{ldr}$ as
-$AR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=1),$ (6)
-$IR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=0),$ (7)
-$M^{sep^{(n)}}_{i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(n)},\\ \rho^{(n)}\times\beta,&if\,(i,j)\in IR^{(n)},\\ 0,&\,otherwise,\end{cases}$ (8)
-$\rho^{(n)}$ $=\frac{N_{AR^{(n)}}}{N_{IR^{(n)}}},$ (9)
-where $\rho^{(n)}$ represents the relative importance of the $IR^{(n)}$ in comparison to $AR^{(n)}$, $\alpha$ and $\beta$ are the balancing parameters between the separated regions, and $N$ is the number of pixels.
-In this way, the distillation region and the $\rho^{(n)}$ value are adaptively adjusted for each training, and distillation is performed with different weights for AR and IR.
-Figure 3: Detailed structure of the proposed CMA module.
-After decomposing the distillation region with AR and IR, the $n$-th aligned features $F^{a^{(n)}}_{rdr}$ are trained to imitate $F^{l}_{ldr}$.
+AFD conducts Activation-aware Feature Matching on the low-level features obtained from both radar and LiDAR branches.
+AFD performs knowledge distillation from the LiDAR BEV features $F^{(l)}_{ldr}$ to the first radar BEV features $F^{(l_{1})}_{rdr}$.
+Simultaneously, it also performs knowledge distillation from $F^{(l)}_{ldr}$ to the second radar BEV features $F^{(l_{2})}_{rdr}$.
+AFD conducts selective knowledge distillation based on regions using active masks.
+We determine the active masks $M^{(l_{1})}_{rdr}$, $M^{(l_{2})}_{rdr}$, and $M^{(l)}_{ldr}$ using the corresponding densified radar features $F^{(l_{1})}_{rdr}$, $F^{(l_{2})}_{rdr}$, and the low-level LiDAR features $F^{(l)}_{ldr}$, respectively, as
+$F^{(l^{\prime})}_{mod,i,j}=\sum^{C}_{c=1}F^{(l^{\prime})}_{mod,c,i,j},$ (4)
+$M^{(l^{\prime})}_{mod,i,j}=\begin{cases}1,&{\rm if}\,F^{(l^{\prime})}_{mod,i,j}>0,\\ 0,&{\rm otherwise},\end{cases}$ (5)
+where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, $i,j,$ and $c$ correspond to height, width, and channel indices, respectively, $mod\in\{ldr,rdr\}$, and $l^{\prime}\in\{l_{1},l_{2},l\}$.
+Using the active masks obtained from (4) and (5), we can identify the active regions (AR) where both radar and LiDAR features are active, and the inactive regions (IR) where radar features are active and LiDAR features are inactive, i.e.,
+$AR^{(l_{n})}$ $=\,(M^{(l_{n})}_{rdr}=1)\,\&\,(M^{(l)}_{ldr}=1),$ (6)
+$IR^{(l_{n})}$ $=\,(M^{(l_{n})}_{rdr}=1)\,\&\,(M^{(l)}_{ldr}=0),$ (7)
+where $n\in[1,2]$.
+AFD minimizes the difference between radar and LiDAR features in both the AR and IR.
+AFD is specifically trained to transform radar features to resemble LiDAR features within AR, while simultaneously learning to suppress radar features within IR.
+Due to the sparsity of LiDAR data, the AR area is typically much smaller than the IR area.
+To prevent the model from overly focusing on knowledge distillation within IR, we scale the loss terms associated with AR and IR based on their relative area ratios.
+We devise adaptive loss weights $W^{(l_{n})}_{sep}$ to balance the loss terms within AR and IR according to their respective pixel counts, i.e.,
+$W^{(l_{n})}_{sep,i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(l_{n})},\\ \rho^{(l_{n})}\times\beta,&if\,(i,j)\in IR^{(l_{n})},\\ 0,&\,otherwise,\end{cases}$ (8)
+where $\rho^{(l_{n})}=\frac{N_{AR^{(l_{n})}}}{N_{IR^{(l_{n})}}}$ represents the relative importance of the $IR^{(l_{n})}$ over $AR^{(l_{n})}$, $\alpha$ and $\beta$ are the intrinsic balancing parameters, and $N_{AR^{(l_{n})}}$ and $N_{IR^{(l_{n})}}$ are the number of pixels in $AR^{(l_{n})}$ and $IR^{(l_{n})}$, respectively.
+The weight associated with IR is determined proportionally by $\rho$, representing the ratio of pixel count for AR to that for IR.
 The distillation loss $L^{(n)}_{low}$ for the $n$-th activation features is given by
-$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{sep^{(n)}}_{i,j}(F^{ldr}_{c,i,j}-F^{a^{(n)}}_{c,i,j})^{2},$ (10)
-where $F^{ldr}$ denotes the LiDAR’s low-level BEV features $F^{l}_{ldr}$.
-The active mask loss is applied between the predicted mask, obtained by applying a sigmoid function to $F^{a(n)}$, and $M^{ldr}$ to ensure the radar’s active mask closely mimics LiDAR.
-The active mask loss $L^{(n)}_{mask}$ is calculated as
-$L^{(n)}_{mask}$ $=\sum^{H}_{i=1}\sum^{W}_{j=1}|M^{ldr}_{i,j}-\sigma(F^{a^{(n)}}_{i,j})|,$ (11)
-where $\sigma(\cdot)$ denotes a sigmoid function.
-We apply AFD twice to the aligned BEV features to more accurately imitate $F^{l}_{ldr}$, and use the average of losses calculated from these multiple AFD as the final AFD loss.
-The final AFD loss function is given by
-$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}(L^{(n)}_{low}+L^{(n)}_{mask}).$ (12)
+$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}W^{(l_{n})}_{sep,i,j}(F^{(l)}_{ldr,c,i,j}-F^{(l_{n})}_{rdr,c,i,j})^{2}.$ (9)
+The final AFD loss $L_{AFD}$ is obtained by averaging the distillation loss functions associated with two densified radar features $F^{(l_{1})}_{rdr}$ and $F^{(l_{2})}_{rdr}$
+$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}L^{(n)}_{low}.$ (10)
 3.4 Proposal-based Feature Distillation
-PFD employs Proposal-level Feature Matching to reduce the discrepancies between the high-level features.
-It utilizes predictions generated from the radar, guiding the radar network to create object features that closely mimic the high-level features of LiDAR, thereby achieving a more accurate alignment in properly identified proposals.
-Additionally, for erroneously identified proposals, such as false positives, PFD directs the model to suppress falsely activated features, thereby mirroring the lower activation levels typically seen in LiDAR features for such instances.
-We employ the radar’s predicted heatmap to adaptively define the crucial region according to radar.
-Additionally, there is a significant scale disparity between the high-level BEV features generated by radar and LiDAR.
-PFD aligns the scales between the two modalities by making the student mimic the normalized values that are applied in the channel direction for both radar and LiDAR.
-DenseEnc perform input $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ by first downsampling to 1/16 of its size and then upsampling to 1/8 to generate the initial high-level BEV feature $F^{h^{(1)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$.
-Subsequently, $F^{h^{(1)}}_{rdr}$ and CMA’s output $F^{a^{(2)}}_{rdr}$ are concatenated and then fed through a convolution layer resulting in the second high-level BEV feature $F^{h^{(2)}}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$.
-Similar to AFD, PFD is applied twice to each of the multiple feature sets $F^{h^{(m)}}_{rdr}$.
-The first step involves thresholding the predicted radar heatmap $H^{cls}_{rdr}$ and the GT heatmap $H^{cls}_{GT}$ to identify the regions of true positives, false positives, and false negatives.
-The regions are identified using the following operations
+PFD employs Proposal-level Feature Matching to transfer knowledge from the high-level features of LiDAR to those of radar.
+This guidance helps the radar network generate object features that closely mimic the high-level features of LiDAR, within object proposals.
+DenseEnc is applied to produce two high-level features $F^{(h_{1})}_{ldr}$ and $F^{(h_{2})}_{ldr}$ from the low-level features $F^{(l)}_{ldr}$ in the LiDAR branch.
+Similarly, DenseEnc also produces two high-level features $F^{(h_{1})}_{rdr}$ and $F^{(h_{2})}_{rdr}$ from $F^{(l_{2})}_{rdr}$ in the radar branch.
+DenseEnc generates the first high-level features $F^{(h_{1})}_{mod}$ using the Conv2D-BN-ReLU block followed by a 2D convolution block and a 2D transposed convolution.
+Here, the Conv2D-BN-ReLU block refers to a sequence of operations comprising 2D Convolution (Conv2D), Batch Normalization (BN), and Rectified Linear Unit (ReLU), while the 2D convolution block consists of six layers of Conv2D-BN-ReLU.
+For LiDAR branch, DenseEnc produces the second high-level LiDAR features $F^{(h_{2})}_{ldr}$ by concatenate $F_{ldr}^{(h_{1})}$ with $F_{ldr}^{(l)}$ and applying the 2D Convolution block again.
+DenseEnc also generates the second high-level radar features $F^{(h_{2})}_{rdr}$ by concatenate $F_{rdr}^{(h_{1})}$ with $F_{rdr}^{(l_{2})}$ and applying the 2D Convolution block in radar branch.
+Finally, the CenterHead is applied to both $F^{(h_{2})}_{ldr}$ and $F^{(h_{2})}_{rdr}$ to produce the classification heatmaps $H^{cls}_{ldr}$ and $H^{cls}_{rdr}$, respectively.
+PFD conducts knowledge distillation solely within the key target regions of high-level BEV features.
+PFD identifies key target regions based on the predicted radar heatmap $H^{cls}_{rdr}$ and the ground truth heatmap $H^{cls}_{GT}$.
+The ground truth heatmap is generated by projecting the 3D centers of ground truth boxes onto the BEV space and applying a Gaussian kernel to represent object locations [44].
+The identified key target regions are labeled as true positives (TP), false positives (FP), or false negatives (FN) according to
 $TP$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}>\sigma),$ (11)
@@
 $FN$ $=\,(H^{cls}_{GT}>\sigma)\,\&\,(H^{cls}_{rdr}<\sigma),$ (13)
-where $H^{cls}_{rdr}$ represents the predicted radar classification heatmap, $H^{cls}_{GT}$ is the GT classification heatmap, and $\sigma$ is the threshold parameter set at 0.1.
-TP and FN are considered important regions according to the foreground, while FP indicates regions where radar is mostly wrong.
-Consequently, TP and FN are categorized into one group, FP into another, leading to the generation of a scaling mask $M^{scale}$ for each defined area.
-The scaling mask is expressed as
-$M^{scale}_{i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (16)
-where $N$ denotes the number of pixels, while $\lambda_{1}$ and $\lambda_{2}$ are balancing parameters for the respective regions.
-The alignment of feature scales between radar and LiDAR involves applying the channel-wise softmax to the high-level BEV features $F^{h}_{ldr}$ and $F^{h^{(m)}}_{rdr}$, which results in $S^{ldr}$ and $S^{h^{(m)}}$, i.e.,
-$S^{mod}_{c,:,:}=\frac{{\text{e}xp}(\frac{F^{mod}_{c,:,:}}{\tau})}{\sum^{C}_{c^{*}=1}{\text{e}xp}(\frac{F^{mod}_{c^{*},:,:}}{\tau})},$ (17)
-where $mod\in\{ldr,h^{(m)}\}$, $\tau$ is a hyper-parameter that controls the smoothing of the distribution and is used as 0.5.
-Similar to AFD, PFD is employed twice to train $S^{h^{(m)}}$ to imitate $S^{ldr}$ and is computed as
-$L^{(m)}_{high}=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{scale}_{i,j}|S^{ldr}_{c,i,j}-S^{h^{(m)}}_{c,i,j}|.$ (18)
-The final PFD loss value is derived from the average of the PFD loss values computed from multiple PFD. The final PFD loss is calculated from
-$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}(L^{(m)}_{high}).$ (19)
-3.5 Overall Loss Function
-We freeze the teacher model that has been pre-trained on LiDAR point clouds and then train the student model end-to-end using radar point clouds.
-Finally, we compute the total loss as the weighted sum of the detection loss and the distillation losses as
+where $\sigma$ represents the threshold parameter set to 0.1 in our setup.
+TP and FN regions encompass areas containing real objects, while FP regions correspond to mis-detected regions.
+Consequently, PFD concentrates on feature alignment in both TP and FN regions while suppressing radar features in FP regions.
+Due to the inherent imbalance in the proportions of these three regions, distinct weights are applied to the distillation loss terms associated with TP, FN, and FP regions.
+The proposal-dependent loss weights $W_{proposal}$ are determined depending on the area of each region
+$W_{proposal,i,j}$ =$=\begin{cases}\frac{\lambda_{1}}{N_{TP}+N_{FN}},&if\,(i,j)\in(TP\cup FN),\\ \frac{\lambda_{2}}{N_{FP}},&if\,(i,j)\in FP,\\ 0,&\,otherwise,\end{cases}$ (14)
+where $\lambda_{1}$ and $\lambda_{2}$ represent the balancing parameters for the respective regions, and $N_{TP}$, $N_{FN}$ and $N_{FP}$ denote the number of pixels within TP, FN, and FP regions, respectively.
+Finally, the distillation loss defined for the $m$th high-level features is weighted by the loss weights $W_{proposal}$ as
+$L^{(m)}_{high}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}W_{proposal,i,j}\left|S^{(h_{m})}_{ldr,c,i,j}-S^{(h_{m})}_{rdr,c,i,j}\right|,$ (15)
+where
+$S^{(h_{m})}_{mod,c,i,j}=\frac{\text{exp}(F^{(h_{m})}_{mod,c,i,j})}{\sum^{C}_{k=1}\text{exp}(F^{(h_{m})}_{mod,k,i,j})},$ (16)
+where $mod\in\{rdr,ldr\}$.
+Both the radar and LiDAR features $F^{(h_{m})}_{rdr}$ and $F^{(h_{m})}_{ldr}$ go through normalization in the channel dimension, resulting in $S^{(h_{m})}_{rdr}$ and $S^{(h_{m})}_{ldr}$, respectively.
+This normalization step aims to align the magnitude scale of the high-level features between the radar and LiDAR branches.
+The final PFD loss $L_{PFD}$ is computed by averaging the distillation loss over $m=1,2$, i.e.,
+$L_{PFD}=\frac{1}{2}\sum^{2}_{m=1}L^{(m)}_{high}.$ (17)
+3.5 Loss Function
+The total loss function used to train RadarDistill is obtained from
 $L_{total}=L_{det}+\gamma L_{AFD}+\delta L_{PFD},$ (18)
-where $\gamma$ and $\delta$ are hyperparameters used to balance the scale of different distillation losses.
+where $\gamma$ and $\delta$ are the regularization parameters for weighting the loss terms $L_{AFD}$ and $L_{PFD}$, respectively.
 4 Experiments
@@
 We conduct the experiments on the nuScenes [1] dataset.
-This dataset contains 700, 150, 150 driving scenes for training, validation, and testing each captured with images from six cameras, points from five radars and one LiDAR.
+This dataset contains 700, 150, 150 driving scenes for training, validation, and testing, respectively.
+The nuScenes dataset gathers radar data using three radar sensors at the vehicle’s front and corners, and two at the rear, providing 360-degree coverage.
+These sensors collectively capture data at a frequency of 13Hz, ensuring broad coverage up to 250 meters and offering velocity measurement accuracy within ±$\pm 0.1km/h$.
 Our evaluation utilizes the official metrics from nuScenes, which include mean average precision (mAP) and the nuScenes detection score (NDS).
@@
   RadarDistill R 20.5 43.7 0.461 0.263 0.525 0.336 0.072
-Table 1: Performance comparison on nuScenes test set.
-The best performed metrics in the radar-only model are bolded.
-”*” denotes models we have reproduced without applying test time augmentation.
+Table 1: Performance evaluation on nuScenes testset.
+The best performed metrics in the radar-only model are marked in bold.
+“*” denotes models we have reproduced without applying test time augmentation.
 Our model achieves the state-of-the-art performance in all metrics.
@@
   RadarDistill R 54.0 15.3 11.3 29.5 5.5 9.2 15.3 0.9 21.7 42.3
-Table 2: Per-class comparison on nuScenes test set.
-The best performed metrics are bolded.
-’C.V’, ’Ped.’, ’Motor.’, and ’T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
+Table 2: Performance evaluation per class on nuScenes testset.
+The best performed metrics are marked in bold.
+‘C.V’, ‘Ped.’, ‘Motor.’, and ‘T.C.’ represent construction vehicle, pedestrian, motorcycle, and traffic cone, respectively.
 Implementation details.
-As a baseline model, we utilized PillarNet-18, which is PillarNet [23] with a ResNet-18 backbone.
-We used the Adam optimizer with a learning rate of 0.001, applying a one-cycle learning rate policy.
-We set weight decay at 0.01, and adjusting momentum between 0.85 and 0.95.
-We trained our baseline models on 4 NVIDIA RTX 3090 GPUs for 20 epochs, employing a total batch size of 16 with the Class-Balanced Grouping and Sampling (CBGS) strategy [44].
-We applied random scene flipping along the axes, random rotation, scaling, translation and ground-truth box sampling as the data augmentation.
-We set our detection range on the $X$ and $Y$ axes to $[-54.0m,54.0m]$ and on the Z-axis to $[-3m,5m]$ , utilizing pillars of $(0.075m,0.075m)$.
-Our proposed model was trained for $40$ epochs and all other training procedures were applied identically to the baseline.
-We adopted the teacher’s weights to initialize the student model, following the inheriting strategy [12].
-In Eq.
-(9), we set the hyper-parameters $\alpha$ to 3e-4 and $\beta$ to 5e-5, while in Eq.
-(16), the hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ were set to $5$ and $1$, respectively.
-Additionally, in Eq.
-(20), we set the parameter $\gamma$ and $\delta$ to $5$ and $25$.
-4.2 Comparison with the State-of-the-Arts
-We evaluated the performance of RadarDistill on nuScenes test set.
-In Table 1, our RadarDistill achieved top-ranking performance on the nuScenes radar-only benchmark, significantly outperforming other methods.
-Specifically, our model achieved a $20.5\%$ mAP and $43.7\%$ NDS, surpassing the previous state-of-the-art (SOTA) method, KPConvPillars [28], by a substantial margin of $+15.6\%$ in mAP and $+29.8\%$ in NDS.
-As shown in Table 2, we compared the performance of our RadarDistill for each class with other camera-only and camera-radar fusion models.
-Our model outperformed the baseline model (PillarNet [23]), especially for the barrier and trailer classes, with gains of at least $+23\%$.
-Furthermore, our model demonstrated the best performance for the car and trailer classes even when compared with other camera-only models and camera-radar fusion model approaches.
-However, the pedestrian and bicycle classes showed only slight performance improvements, largely due to the inherent limitations of radar sensors, as these objects are smaller and have few radar points in their ground truth (GT) boxes.
+As our baseline model, we employed PillarNet-18, which corresponds to PillarNet [23] with a ResNet-18 backbone.
+We utilized the Adam optimizer with a learning rate of 0.001, implementing a one-cycle learning rate policy.
+The weight decay was set to 0.01 with the momentum scheduled from 0.85 to 0.95.
+We trained the PillarNet-18 model for 20 epochs with a batch size of 16.
+We used the Class-Balanced Grouping and Sampling (CBGS) strategy [45] to mitigate class imbalance issue.
+Data augmentation techniques were applied, including random scene flipping along the $X$ and $Y$ axes, random rotation, scaling, translation, and ground-truth box sampling.
+Our detection range was set to $[-54m,54m]$ for both the $X$ and $Y$ axes, and $[-5m,3m]$ for the $Z$ axis.
+We used pillars of dimensions $(0.075m,0.075m)$.
+Our proposed RadarDistill model was trained for 40 epochs.
+The rest of training setup follows the same training setup as the baseline model.
+We initialized the radar backbone network using the pre-trained LiDAR backbone network, following the Inheriting Strategy presented in [12].
+The hyperparameters $\alpha$ and $\beta$ in (8) were set to $3\times 10^{-4}$ and $5\times 10^{-5}$, respectively.
+Additionally, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ in (14) were set to $5$ and $1$, respectively.
+The parameters $\gamma$ and $\delta$ in (18) were set to $5$ and $25$, respectively.
+The entire training was conducted on 4 NVIDIA RTX 3090 GPUs.
+4.2 Performance Comparison
+Table 1 presents the performance of RadarDistill on nuScenes testset.
+Our RadarDistill achieves significant performance improvements compared to other radar-based object detectors.
+Specifically, RadarDistill achieves a mAP of 20.5% and a NDS of 43.7%, outperforming the previous state-of-the-art (SOTA) method, KPConvPillars [29], by a considerable margin of +15.6% in mAP and +29.8% in NDS.
+Table 2 displays the AP performance of RadarDistill for each class.
+We compare our RadarDistill with other camera-only and camera-radar fusion-based object detectors.
+RadarDistill exhibits superior performance compared to the baseline, PillarNet-18 [23] across all classes.
+Notably, it achieves performance gains of +12.2% and +23% for the Car and Trailer classes, respectively.
+Furthermore, RadarDistill outperforms other camera-only and sensor fusion models in the Car class, which is surprising given the limited resolution of radar data compared to camera data.
 4.3 Ablation Studies
-We conducted several ablation studies using the nuScenes validation set.
-We trained our model on 1/7 of training set and used the complete validation set for evaluation to reduce the time spent on experiments.
+We conducted ablation studies on the nuScenes validation set to assess the impact of each proposed idea.
+Our model was trained using $1/7$ of the training set and evaluated on the entire validation set.
 Component Analysis.
-We present an ablation study to evaluate the impact of each component of our proposed model, as shown in Table 3.
-In #2, when CMA is applied to the baseline, there’s an overall enhancement in performance; yet, the improvements are relatively low.
-Comparing #2 with #3 and #4, where CMA is combined with AFD and PFD respectively, we see a significant performance gain.
-For #3, there is an increase in mAP by $+4.5\%$ and NDS by $+4.4\%$.
-In the case of #4, mAP improves by $+1.8\%$ and NDS by $+3.0\%$.
-Additionally, when all components are used together as in #6, we can see that it achieves the best performance across all metrics.
-Lastly, to understand the influence of CMA in the distillation process, #5 was tested without it.
-The results show a significant difference in performance, with an increase of $+4.2\%$ in mAP and $+4.0\%$ in NDS when CMA is included, demonstrating that the CMA structure performs a critical role in our model.
-Comparative Analysis of Distillation Regions. We conducted experiments to analyze the importance of distillation regions during the knowledge distillation (KD) process.
-Table 4 presents an ablation study conducted to evaluate the impact of selecting different regions when applying low-level feature distillation.
-When distillation is performed on the entire area without separate regions, we observe a performance improvement of $+4.2\%$ in Car AP, $+2.2\%$ in mAP, and $+2.9\%$ in NDS.
-Applying distillation to the gaussian heatmap areas centered around the GT boxes led to an increase in Car AP and mAP compared to the total, but a decrease in NDS.
-Dividing the distillation between foreground and background regions, compared to the total, resulted in an increase in both Car AP and mAP, but a slight decrease in NDS.
-Conversely, dividing the feature distillation by our proposed activation regions yielded the best performance across all metrics, with improvements of $+6.8\%$ in Car AP, $+2.3\%$ in mAP, and $+1.5\%$ in NDS compared to total.
-Next, Table 5 presents an experiment comparing the performance between previous cross-modality KD methods and PFD’s approach in applying high-level feature distillation.
-Here, the first row is based on the application of AFD according to our proposed method.
-When distillation is performed on the entire area, we observe a slight decrease in Car AP and mAP by $-0.6\%$ and $-0.1\%$ respectively, while NDS shows a marginal increase of $0.1\%$, indicating a minimal change in performance.
-Even when comparing the gaussian heatmap areas centered around the GT boxes, only Car AP shows a slight increase $0.1\%$ with the rest being identical to the total.
-Furthermore, applying KD based on LiDAR prediction results also did not yield any notable performance improvements.
-In contrast, our method, which applies KD based on radar prediction boxes, showed an improvement of over $0.4\%$ across all metrics compared to the total, demonstrating the best performance.
-Scale Alignment in PFD.
-We conducted experiments to examine the impact of aligning the scale between teacher and student in PFD process in Table 6.
-When applying PFD without normalization, we observe a substantial decrease of $-5.9\%$ in Car AP, $-2.4\%$ in mAP, and $-1.9\%$ in NDS compared to before applying PFD.
-Conversely, after implementing normalization and PFD, we observe an increase of $+0.3\%$ in Car AP, $+0.3\%$ in mAP, and $+1.0\%$ in NDS.
-Apply radar-camera fusion model.
-Table 7 shows a comparison of the performance when using radar features enhanced by our method in the radar-camera fusion model.
-The enhanced radar and camera features were fused using the BEVFusion [20] method, with its detailed structure and training method provided in the Supplementary Materials.
-We observe a significant increase of $+1.8\%$ in Car AP, $+1.3\%$ in mAP, and $+1.1\%$ in NDS.
+Table 3 illustrates the performance enhancements achieved by each component of our proposed model.
+Integrating CMA into the PillarNet-18 baseline results in a 2% improvement in NDS performance.
+With both CMA and AFD enabled, the NDS performance further improves by 4.4%.
+Activating PFD yields a 1.0% additional increase in NDS performance.
+We also tried disabling CMA while other components are enabled in RadarDistill.
+In this case, the NDS performance drops by 4%, highlighting the critical role of CMA in our RadarDistill model.
+Proposed Distillation Method. We compare the proposed Distillation method with other well known KD methods. We consider the following methods
+  • Baseline*: This baseline is constructed by integrating CMA into the PillarNet-18 baseline.
+  • Baseline**: This baseline is constructed by applying CMA and AFD to the PillarNet-18 baseline.
+  • Complete [22]: This method applies the fixed weight across entire areas of the low-level BEV features.
+  • Gaussian [26]: This approach generates a 2D Gaussian mask based on the centers of GT boxes and assigns higher weighted distillation loss to areas closer to the foreground.
+  • FG/BG [7]: This strategy separates the foreground and background using GT boxes, applying different weighted distillation losses for each.
+Table 4 evaluates the performance of AFD in comparison with other KD methods.
+We applied AFD and these methods to the Baseline*, where the low-level BEV features are produced by CMA.
+We observe that the proposed AFD achieves significantly better performance than other KD methods.
+Particularly, it achieves the performance gains of 3.7% in AP in car category, 1.3% in mAP, and 2.0% in NDS over the FG/BG method.
+Table 5 compares PFD with other KD methods when applied to high-level BEV features.
+We applied these methods to Baseline** where the high-level features are produced by CMA and AFD.
+We also confirm that the proposed PFD achieves higher detection performance gains than other KD methods.
+Specifically, it yields the performance improvements of 0.5% in AP for the car category, 0.2% in mAP, and 1.0% in NDS over the FG/BG method.
 […]
-1 5.4 27.3
-2 6.4 29.3
-3 10.9 33.7
-4 8.2 32.3
-5 7.0 30.7
-6 11.2 34.7
-Table 3: Ablation study of components in Radardistill. “CMA”: Cross-Modality Alignment. “AFD”: Activation-based Feature Distillation. “PFD”: Proposal-based Feature Distillation.
+5.4 27.3
+6.4 29.3
+10.9 33.7
+11.2 34.7
+7.0 30.7
+Table 3: Ablation study for evaluating the contribution from each component of Radardistill on nuScenes validation set.
 Region AFD
 […]
-W/O AFD 34.8 6.4 29.3
-Total 39.0 8.6 32.2
-GT 40.4 8.8 29.5
-FG / BG 42.1 9.6 31.7
-Activation 45.8 10.9 33.7
-Table 4: Ablation study of distillation regions in AFD. “W/O AFD”: Applying only CMA without distillation. “FG/BG”: Foreground and background region.
+Baseline* 34.8 6.4 29.3
+Complete [22] 39.0 8.6 32.2
+Gaussian [26] 40.4 8.8 29.5
+FG/BG [7] 42.1 9.6 31.7
+Our AFD 45.8 10.9 33.7
+Table 4: Comparison of AFD with different KD methods evaluated on nuScenes validation set.
+Impact of Scale Normalization in PFD.
+Next, we investigate the impact of scale normalization in PFD.
+Table 6 presents a comparison of performances with and without scale normalization.
+We observe that our scale normalization results in a significant improvement of 6.2% in AP (CAR category), 2.7% in mAP, and 2.9% in NDS.
+This underscores the substantial difference in magnitude scale between radar and LiDAR high-level features, and demonstrates that narrowing this gap significantly boosts the effect of knowledge distillation.
+Comparison when applied to Radar-Camera Fusion.
+Next, we investigate whether the improved features generated by RadarDistill can also lead to performance enhancements in radar-camera fusion methods.
+Table 7 presents the detection performance when the proposed RadarDistill is intergrated into the radar-camera fusion method, BEVFusion [20].
+Because BEVFusion was originally designed for LiDAR-camera fusion, we adapted its design for radar-camera fusion following [15].
+Then, RadarDistill was simply incorporated to the radar encoding network in BEVFusion and the entire model was trained end to end.
+We note that RadarDistill yields performance improvements of 1.8% in Car AP, 1.3% in mAP, and 1.1% in NDS over the baseline BEVFusion model.
+We believe that further enhancements in performance could be achieved by applying more sophisticated fusion strategies, a direction we leave for future research.
 Region PFD
 […]
-W/O PFD 45.8 10.9 33.7
-Total 45.2 10.8 33.8
-GT 45.9 10.8 33.8
-LiDAR prediction 45.4 10.8 34.0
-Radar prediction 46.1 11.2 34.7
-Table 5: Ablation study of distillation regions in PFD.
-“W/O PFD”: Applying only CMA and AFD.
-“LiDAR prediction” : Using only LiDAR prediction results.
-“Radar prediction” : Using only Radar prediction results
+Baseline** 45.8 10.9 33.7
+Complete [22] 45.2 10.8 33.8
+Gaussian [26] 45.9 10.8 33.8
+FG/BG [7] 45.6 11.0 33.7
+Our PFD 46.1 11.2 34.7
+Table 5: Comparison of PFD with different KD methods evaluated on nuScenes validation set.
 […]
-W/O PFD 45.8 10.9 33.7
-W/O Norm 39.9 8.5 31.8
-Norm 46.1 11.2 34.7
-Table 6: Ablation study of scale alignment in PFD. “W/O PFD”: Applying only CMA and AFD. “W/O Norm”: Applying PFD without normalizing. “Norm”: Applying normalizing and PFD.
+PFD w/o norm 39.9 8.5 31.8
+PFD 46.1 11.2 34.7
+Table 6: Ablation study for evaluating the impact of scale normalization used in PFD on nuScenes validation set. PFD w/o norm: PFD without scale normalization. PFD: PFD with scale normalization.
 […]
@@
 RadarDistill-CR C,R 67.7 39.6 46.4
-Table 7: Ablation study of applying to a radar-camera fusion model.
-”*” denotes models we have reproduced without applying test time augmentation.
-Our model achieves slight performance enhancements in all metrics.
-5 Conclusion
-In our study, we introduced a new radar-based 3D object detection method, RadarDistill, which leveraged LiDAR data via knowledge distillation for enhanced performance.
-Our approach successfully guided the radar encoding network to generate radar features that closely resemble the detailed and semantically rich features of LiDAR.
-CMA enhanced the student network’s expressive power beyond that of the teacher network, allowing the student network to learn the complex distribution of LiDAR features more effectively.
-Through Activation-aware Feature Matching and Proposal-level Selective Feature Matching, both AFD and PFD successfully reduced the discrepancy from LiDAR feature distributions, particularly in important regions within the intermediate features.
-Our experiments demonstrated that RadarDistill achieved the state-of-the-art performance in radar-only object detection and offered significant performance gains in camera-radar fusion scenarios.
-We leave the design of more sophisticated KD framework specifically tailored for camera-radar fusion for future work.
+Table 7: Ablation study for evaluating the impact of RadarDistill on radar-camera fusion models. “*” denotes models we have reproduced without applying test time augmentation.
+5 Conclusions
+In this study, we introduced RadarDistill, a novel radar-based 3D object detection method aimed at enhancing radar data features by leveraging information from LiDAR data through knowledge distillation.
+We proposed three effective techniques—CMA, AFD, and PFD—to make the challenging task of transferring knowledge from LiDAR to radar successful.
+Our approach emphasizes guiding the radar encoding network to generate features that closely resemble the semantically rich features of LiDAR.
+CMA densifies radar features, aiding the radar encoding network in learning the complex distribution of LiDAR features better.
+AFD and PFD target the reduction of discrepancies between radar and LiDAR features in both low-level and high-level BEV features.
+Our evaluation demonstrated that RadarDistill can yield significant performance improvements over both radar-based and radar-camera fusion-based baselines, establishing state-of-the-art performance in radar-only object detection.
+6 Acknowledgement
+This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.2021-0-01343-004, Artificial Intelligence Graduate School Program (Seoul National University)] and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No.2020R1A2C2012146].